Find the nth partial sum of the arithmetic sequence. an 5n + 2; n = 10 S10 =
Answer #1

Solution! Given that- an = 5m +2 = 10 a = 5X1+& = 5+2 = 7 and ago 5X 10 +2 = 50 +2 = 52 We know that the formula of nth parti

If any doubt please comment